# Beyond the Main Sequence: Binary Evolution Pathways to Blue Stragglers in the Gaia Era

### v1

@@ Acknowledgements.
P.K.N. and T.H.P. acknowledge support by the ANID BASAL projects FB210003. T.H.P. gratefully acknowledges support through FONDECYT Regular Project No. 1201016.cluster cluster RA Dec log(age/yr) Av distance [Fe/H] Teff,IRFM Teff,isochrone NBSS NMS fBSS/MS name type (deg) (deg) (mag) (pc) (dex) (K) (K) (1) (2) (3) (4) (5) (6) (7) (8) (9) (10) (11) (12) (13) NGC104 GC 6.175 72.024 10.0700.011 0.124 4367 0.65 590392 5995 415 13691 0.060.002 NGC288 GC 13.381 26.529 10.0640.002 0.031 9000 0.8 6066116 6173 150 – – NGC362 GC 15.885 70.781 10.0680.013 0.155 9090 1.3 6248160 6368 137 1578 0.1710.011 ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ Berkeley39 OC 116.893 4.735 9.811 0.029 0.394 4156 0.0 5932 84 6038 22 264 0.14 0.025 Berkeley32 OC 104.738 6.379 9.686 0.022 0.553 3321 0.3 6286 125 6403 14 139 0.151 0.035 Collinder261 OC 189.802 68.492 9.898 0.065 0.999 2706 0.1 5772 72 5865 64 460 0.228 0.025 ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ 555The list of columns represented here is: (1) Name of the cluster; (2) whether it is a globular cluster (GC) or open cluster (OC); (3) and (4) represents central coordinates of the cluster, (5) age of the cluster in log scale, (6) extinction (AV) towards the cluster, (7) distance to the cluster, (8) metallicity of the cluster, (9) mean effective temperature of BSS populations in that cluster using IRFM method, (10) mean effective temperature of BSS populations in that cluster using isochrone method; (11) number of BSSs detected in that cluster; (12) number of main sequence (MS) stars in the luminosity range between the MS turn-off (MSTO) and MSTO+1 mag in that cluster; (13) fractional BSSs population to MS populations.
Isochrone Method cluster type Gaia source id Teff,IRFM Teff log(age/yr) M log(L/L⊙) logg Teff,xpp loggxpp Me Mesequence (K) (K) (M⊙) (K) (mag) (mag) (1) (2) (3) (4) (5) (6) (7) (8) (9) (10) (11) (12) (13) (14) (15) NGC2141 OC 3341815582307132416 636382 651059 9.2220.015 1.700.01 1.0400.006 3.8330.010 – – – – 0.17 lower 2.237 0.597 NGC2141 OC 3341815582307133056 639184 646479 9.2220.015 1.720.00 1.0700.003 3.8000.020 – – – – 0.19 lower 2.176 0.589 NGC2141 OC 3341815681086838272 6736101 685595 9.0870.077 1.560.01 0.8000.005 4.1220.025 – – – – 0.08 lower 2.807 0.500 ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ NGC104 GC 4689572883825672192 6422126 656746 9.7400.027 1.010.01 0.5500.004 4.1190.015 6831136 4.1420.123 0.28 upper 3.538 0.566 NGC104 GC 4689573433581464192 6262109 638426 9.8420.024 0.930.01 0.3500.003 4.2260.009 6896166 4.2160.2 0.18 lower 4.043 0.609 NGC104 GC 4689574086416430464 6669154 683835 9.5050.035 1.080.01 0.4900.002 4.2710.009 6652159 4.1950.142 0.37 upper 3.651 0.505 ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ 666The list of columns presented here is (1) the name of the cluster to which the BSS belong; (2) whether the cluster is classified as an open cluster (OC) or globular cluster (GC); (3) source id of the BSS as per the Gaia DR3; (4) effective temperature of the BSS based on IRFM method; (5) – (9) effective temperature, age, mass, luminosity and surface gravity of the BSS, respectively, estimated using isochrone method; (10) – (11) effective temperature and of the BSS estimated from Gaia-XPP spectrum. (12) fractional mass excess of the BSS based compared to MSTO mass; (13) location of the BSS based on mass excess, as mentioned in Figure 9; (14)-(15) BP magnitude and (BP-RP) colour in absolute scale.
@@ Appendix A Validation of our membership criteria:
To validate our BSS membership criteria as well as all other cluster members, we compare our cluster members with the literature. We only compare the clusters which show the presence of BSSs as we only focus on these clusters in our further analyses. The most recent and extensive membership catalogs using Gaia data are given by Vasiliev & Baumgardt (2021); Hunt & Reffert (2023) for GCs and Cantat-Gaudin et al. (2020a); Hunt & Reffert (2023) for OCs. Vasiliev & Baumgardt (2021) provided membership probability of all stars in 170 GCs using the Gaussian mixture model on Gaia-EDR3 data, where stars with a membership probability greater than 90% are considered to be cluster members. We find more than 90% match in membership for all 41 GCs, with a median crossmatch accuracy of 99.7%. All BSSs detected in our GCs sample are also classified as cluster members in Vasiliev & Baumgardt (2021). Hunt & Reffert (2023) used HDBSCAN algorithm on Gaia DR3 data down to magnitude G 20 mags to recover 7167 clusters, including 134 globular clusters. For GCs, we find that the crossmatch between our catalog and Hunt & Reffert (2023) varies from 15% (for NGC 6656; very close to the Galactic plane) to 95% based on the distance of the clusters from the Galactic plane, with a median crossmatch accuracy of 84%. NGC 104 is not studied by Hunt & Reffert (2023), hence we are unable to make any crossmatch with this cluster. The membership method by Hunt & Reffert (2023) is found to have a bias on whether the cluster located in a crowded or sparse field, which is not found in the methods of Vasiliev & Baumgardt (2021). In the top panel of Figure 12, we show CMDs of two GCs (NGC 3201 and NGC 288) to demonstrate the comparison of the cluster membership obtained from our method (blue), Vasiliev & Baumgardt (2021) in orange and Hunt & Reffert (2023) in green. The figure shows a good agreement between all three methods.
@@
We find a better crossmatch with Hunt & Reffert (2023) for OCs. We find a crossmatch of more than 50% for all cluster (except UBC252 with 19%) and with median crossmatch value of 94%. The unmatched sources are mainly fainter than 18 mags in G band and more importantly, all BSSs in our sample are also classified as cluster members in Hunt & Reffert (2023). Dias et al. (2021) used the classic maximum likelihood approach (Dias et al. 2014; Monteiro et al. 2020) on Gaia-DR2 data to study 2000 open clusters for stars brighter than 18 mags in the G band. The comparison shows that the crossmatch fraction varies from 60% to 96% with a median value of 88%. The difference in the crossmatch is mainly found towards the fainter G magnitudes. When cross-matched only with our BSS sample, we find more than 97% are also classified as members in Dias et al. (2021). Cantat-Gaudin et al. (2020a) applied an artificial neural network in Gaia DR2 photometry down to G = 18 mags to estimate the cluster membership of 2000 OCs. Comparing our cluster members with stars brighter than 18 mags in G band, we find that crossmatch fraction varies from cluster to cluster with a range from 40% to 95%. There are mainly two reasons behind the differences in crossmatch fraction: Cantat-Gaudin et al. (2020a) used Gaia DR2 for the membership instead of DR3 with better astrometry and was unable to recover the cluster member properly towards the fainter magnitude limit. For example, there are no cluster member present with G17 mags for the cluster UBC 252, which leads to only 41% of our cluster member. While we consider the match for BSSs, we notice that out of 434 BSSs candidate in our catalog, 409 (94%) are also part of the cluster member in Cantat-Gaudin et al. (2020a). In the bottom panel of Figure 12, we present CMDs of two OCs (NGC 188 and Berkeley 39) as examples to display the comparison between our cluster membership (open blue box) to those obtained by Hunt & Reffert (2023) marked in orange points, Cantat-Gaudin et al. (2020a) marked in green cross, and Dias et al. (2021) in red points. The cluster memberships are found to be well-matched among the different methods. Jadhav & Subramaniam (2021) also provides an extensive catalog of BSSs candidates in OCs. However, they used the primary catalog of OCs by Cantat-Gaudin et al. (2020a). So, we did not perform any details crossmatch with Jadhav & Subramaniam (2021). Therefore, we find that our cluster membership matches well with the literature and almost all BSSs are also classified as members in the literature, which also validates the membership determination algorithm used in our study.

### v2

@@ Acknowledgements.
P.K.N. and T.H.P. acknowledge support by the ANID BASAL projects FB210003. T.H.P. gratefully acknowledges support through the FONDECYT Regular Project No. 1201016.References
- Andrae et al. (2023) Andrae, R., Fouesneau, M., Sordo, R., et al. 2023, A&A, 674, A27
- Astropy Collaboration et al. (2013) Astropy Collaboration, Robitaille, T. P., Tollerud, E. J., et al. 2013, A&A, 558, A33
- Baldwin et al. (2016) Baldwin, A. T., Watkins, L. L., van der Marel, R. P., et al. 2016, ApJ, 827, 12
- Baumgardt et al. (2023) Baumgardt, H., Hénault-Brunet, V., Dickson, N., & Sollima, A. 2023, MNRAS, 521, 3991
- Baumgardt & Vasiliev (2021) Baumgardt, H. & Vasiliev, E. 2021, MNRAS, 505, 5957
- Beccari et al. (2019) Beccari, G., Ferraro, F. R., Dalessandro, E., et al. 2019, ApJ, 876, 87
- Bisht et al. (2022) Bisht, D., Zhu, Q., Elsanhoury, W. H., et al. 2022, AJ, 164, 171
- Bressan et al. (2012) Bressan, A., Marigo, P., Girardi, L., et al. 2012, MNRAS, 427, 127
- Burbidge & Sandage (1958) Burbidge, E. M. & Sandage, A. 1958, ApJ, 128, 174
- Cadelano et al. (2022) Cadelano, M., Ferraro, F. R., Dalessandro, E., et al. 2022, ApJ, 941, 69
- Cantat-Gaudin et al. (2020a) Cantat-Gaudin, T., Anders, F., Castro-Ginard, A., et al. 2020a, A&A, 640, A1
- Cantat-Gaudin et al. (2020b) Cantat-Gaudin, T., Anders, F., Castro-Ginard, A., et al. 2020b, A&A, 640, A1
- Cantat-Gaudin et al. (2018) Cantat-Gaudin, T., Jordi, C., Vallenari, A., et al. 2018, A&A, 618, A93
- Cardelli et al. (1989) Cardelli, J. A., Clayton, G. C., & Mathis, J. S. 1989, ApJ, 345, 245
- Carretta et al. (2009) Carretta, E., Bragaglia, A., Gratton, R., D’Orazi, V., & Lucatello, S. 2009, A&A, 508, 695
- Chen & Han (2009) Chen, X. & Han, Z. 2009, MNRAS, 395, 1822
- Clarkson et al. (2011) Clarkson, W. I., Sahu, K. C., Anderson, J., et al. 2011, ApJ, 735, 37
- Cohen (1988) Cohen, J. 1988, Statistical Power Analysis for the Behavioral Sciences, 2nd edn. (Hillsdale, NJ: Routledge)
- Contreras Ramos et al. (2012) Contreras Ramos, R., Ferraro, F. R., Dalessandro, E., Lanzoni, B., & Rood, R. T. 2012, ApJ, 748, 91
- Cordoni et al. (2018) Cordoni, G., Milone, A. P., Marino, A. F., et al. 2018, ApJ, 869, 139
- Cordoni et al. (2023) Cordoni, G., Milone, A. P., Marino, A. F., et al. 2023, A&A, 672, A29
- Cordoni et al. (2020) Cordoni, G., Milone, A. P., Mastrobuono-Battisti, A., et al. 2020, ApJ, 889, 18
- Culpan et al. (2021) Culpan, R., Pelisoli, I., & Geier, S. 2021, A&A, 654, A107
- Curran (2014) Curran, P. A. 2014, arXiv e-prints, arXiv:1411.3816
- Da Costa (1984) Da Costa, G. S. 1984, ApJ, 285, 483
- Dalessandro et al. (2013) Dalessandro, E., Ferraro, F. R., Massari, D., et al. 2013, ApJ, 778, 135
- Dattatrey et al. (2023) Dattatrey, A. K., Yadav, R. K. S., Rani, S., et al. 2023, ApJ, 943, 130
- De Angeli et al. (2023) De Angeli, F., Weiler, M., Montegriffo, P., et al. 2023, A&A, 674, A2
- De Marco et al. (2005) De Marco, O., Shara, M. M., Zurek, D., et al. 2005, ApJ, 632, 894
- Dias et al. (2014) Dias, W. S., Monteiro, H., Caetano, T. C., et al. 2014, A&A, 564, A79
- Dias et al. (2021) Dias, W. S., Monteiro, H., Moitinho, A., et al. 2021, MNRAS, 504, 356
- Evans et al. (2018) Evans, D. W., Riello, M., De Angeli, F., et al. 2018, A&A, 616, A4
- Ferraro et al. (2009) Ferraro, F. R., Beccari, G., Dalessandro, E., et al. 2009, Nature, 462, 1028
- Ferraro et al. (2012) Ferraro, F. R., Lanzoni, B., Dalessandro, E., et al. 2012, Nature, 492, 393
- Ferraro et al. (2018) Ferraro, F. R., Lanzoni, B., Raso, S., et al. 2018, ApJ, 860, 36
- Ferraro et al. (2023) Ferraro, F. R., Lanzoni, B., Vesperini, E., et al. 2023, ApJ, 950, 145
- Ferraro et al. (1997) Ferraro, F. R., Paltrinieri, B., Fusi Pecci, F., et al. 1997, A&A, 324, 915
- Fiorentino et al. (2014) Fiorentino, G., Lanzoni, B., Dalessandro, E., et al. 2014, ApJ, 783, 34
- Gaia Collaboration et al. (2018) Gaia Collaboration, Brown, A. G. A., Vallenari, A., et al. 2018, A&A, 616, A1
- Gaia Collaboration et al. (2016) Gaia Collaboration, Prusti, T., de Bruijne, J. H. J., et al. 2016, A&A, 595, A1
- Gaia Collaboration et al. (2023) Gaia Collaboration, Vallenari, A., Brown, A. G. A., et al. 2023, A&A, 674, A1
- González Hernández & Bonifacio (2009) González Hernández, J. I. & Bonifacio, P. 2009, A&A, 497, 497
- Gosnell et al. (2015) Gosnell, N. M., Mathieu, R. D., Geller, A. M., et al. 2015, ApJ, 814, 163
- Harris (2010) Harris, W. E. 2010, arXiv e-prints, arXiv:1012.3224
- Hills & Day (1976) Hills, J. G. & Day, C. A. 1976, Astrophys. Lett., 17, 87
- Hunt & Reffert (2023) Hunt, E. L. & Reffert, S. 2023, A&A, 673, A114
- Hurley et al. (2002) Hurley, J. R., Tout, C. A., & Pols, O. R. 2002, MNRAS, 329, 897
- Iben (2013) Iben, Icko, J. 2013, Stellar Evolution Physics, Volume 2: Advanced Evolution of Single Stars
- Jadhav et al. (2019) Jadhav, V. V., Sindhu, N., & Subramaniam, A. 2019, ApJ, 886, 13
- Jadhav & Subramaniam (2021) Jadhav, V. V. & Subramaniam, A. 2021, MNRAS, 507, 1699
- Jiang (2022) Jiang, D. 2022, ApJ, 940, 97
- Jofré et al. (2016) Jofré, P., Jorissen, A., Van Eck, S., et al. 2016, A&A, 595, A60
- Knigge et al. (2009) Knigge, C., Leigh, N., & Sills, A. 2009, Nature, 457, 288
- Leigh et al. (2011) Leigh, N., Sills, A., & Knigge, C. 2011, MNRAS, 416, 1410
- Leiner & Geller (2021) Leiner, E. M. & Geller, A. 2021, ApJ, 908, 229
- Li et al. (2023) Li, C., Zhong, J., Qin, S., & Chen, L. 2023, A&A, 672, A81
- Lombardi et al. (2002) Lombardi, James C., J., Warren, J. S., Rasio, F. A., Sills, A., & Warren, A. R. 2002, ApJ, 568, 939
- Mapelli et al. (2009) Mapelli, M., Ripamonti, E., Battaglia, G., et al. 2009, MNRAS, 396, 1771
- Mapelli et al. (2006) Mapelli, M., Sigurdsson, S., Ferraro, F. R., et al. 2006, MNRAS, 373, 361
- Marigo et al. (2008) Marigo, P., Girardi, L., Bressan, A., et al. 2008, A&A, 482, 883
- Mathys (1987) Mathys, G. 1987, A&AS, 71, 201
- McCrea (1964) McCrea, W. H. 1964, MNRAS, 128, 147
- Montegriffo et al. (2023) Montegriffo, P., De Angeli, F., Andrae, R., et al. 2023, A&A, 674, A3
- Monteiro et al. (2020) Monteiro, H., Dias, W. S., Moitinho, A., et al. 2020, MNRAS, 499, 1874
- Mucciarelli et al. (2021) Mucciarelli, A., Bellazzini, M., & Massari, D. 2021, A&A, 653, A90
- Naoz & Fabrycky (2014) Naoz, S. & Fabrycky, D. C. 2014, ApJ, 793, 137
- Nardiello et al. (2018) Nardiello, D., Libralato, M., Piotto, G., et al. 2018, MNRAS, 481, 3382
- Overbeek et al. (2017) Overbeek, J. C., Friel, E. D., Donati, P., et al. 2017, A&A, 598, A68
- Pandey et al. (2021) Pandey, S., Subramaniam, A., & Jadhav, V. V. 2021, MNRAS, 507, 2373
- Perets & Fabrycky (2009) Perets, H. B. & Fabrycky, D. C. 2009, ApJ, 697, 1048
- Piotto et al. (2015) Piotto, G., Milone, A. P., Bedin, L. R., et al. 2015, AJ, 149, 91
- Preston & Sneden (2000) Preston, G. W. & Sneden, C. 2000, AJ, 120, 1014
- Privon et al. (2020) Privon, G. C., Ricci, C., Aalto, S., et al. 2020, ApJ, 893, 149
- Rain et al. (2021) Rain, M. J., Ahumada, J. A., & Carraro, G. 2021, A&A, 650, A67
- Rao et al. (2023a) Rao, K. K., Bhattacharya, S., Vaidya, K., & Agarwal, M. 2023a, MNRAS, 518, L7
- Rao et al. (2023b) Rao, K. K., Vaidya, K., Agarwal, M., Balan, S., & Bhattacharya, S. 2023b, MNRAS, 526, 1057
- Rao et al. (2021) Rao, K. K., Vaidya, K., Agarwal, M., & Bhattacharya, S. 2021, MNRAS, 508, 4919
- Raso et al. (2017) Raso, S., Ferraro, F. R., Dalessandro, E., et al. 2017, ApJ, 839, 64
- Rodrigo & Solano (2020) Rodrigo, C. & Solano, E. 2020, in XIV.0 Scientific Meeting (virtual) of the Spanish Astronomical Society, 182
- Rodrigo et al. (2012) Rodrigo, C., Solano, E., & Bayo, A. 2012, SVO Filter Profile Service Version 1.0, IVOA Working Draft 15 October 2012
- Sandage (1953) Sandage, A. R. 1953, AJ, 58, 61
- Schultz & Wiemer (1975) Schultz, G. V. & Wiemer, W. 1975, A&A, 43, 133
- Shao & Li (2016) Shao, Y. & Li, X.-D. 2016, ApJ, 833, 108
- Shara et al. (1997) Shara, M. M., Saffer, R. A., & Livio, M. 1997, ApJ, 489, L59
- Sills et al. (2009) Sills, A., Karakas, A., & Lattanzio, J. 2009, ApJ, 692, 1411
- Simunovic (2017) Simunovic, M. 2017, PhD thesis, Ruprecht-Karls University of Heidelberg, Germany
- Simunovic & Puzia (2016) Simunovic, M. & Puzia, T. H. 2016, MNRAS, 462, 3401
- Simunovic et al. (2014) Simunovic, M., Puzia, T. H., & Sills, A. 2014, ApJ, 795, L10
- Sindhu et al. (2019) Sindhu, N., Subramaniam, A., Jadhav, V. V., et al. 2019, ApJ, 882, 43
- Stȩpień & Kiraga (2015) Stȩpień, K. & Kiraga, M. 2015, A&A, 577, A117
- Subramaniam et al. (2016) Subramaniam, A., Sindhu, N., Tandon, S. N., et al. 2016, ApJ, 833, L27
- Vaidya et al. (2022) Vaidya, K., Panthi, A., Agarwal, M., et al. 2022, MNRAS, 511, 2274
- Vasiliev & Baumgardt (2021) Vasiliev, E. & Baumgardt, H. 2021, MNRAS, 505, 5978
- Virtanen et al. (2020) Virtanen, P., Gommers, R., Oliphant, T. E., et al. 2020, Nature Methods, 17, 261
- Wasserstein & Lazar (2016) Wasserstein, R. L. & Lazar, N. A. 2016, The American Statistician, 70, 129
- Zhang et al. (2023) Zhang, X., Green, G. M., & Rix, H.-W. 2023, MNRAS, 524, 1855
@@ Appendix A Validation of our membership criteria:
To validate our BSS membership criteria as well as all other cluster members, we compare our cluster members with the literature. We only compare the clusters which show the presence of BSSs as we only focus on these clusters in our further analyses. The most recent and extensive membership catalogs using Gaia data are given by Vasiliev & Baumgardt (2021); Hunt & Reffert (2023) for GCs and Cantat-Gaudin et al. (2020a); Hunt & Reffert (2023) for OCs. Vasiliev & Baumgardt (2021) provided membership probability of all stars in 170 GCs using the Gaussian mixture model on Gaia-EDR3 data, where stars with a membership probability greater than 90% are considered to be cluster members. We find more than 90% matches in membership for all 41 GCs, with a median crossmatch accuracy of 99.7%. All BSSs detected in our GCs sample are also classified as cluster members in Vasiliev & Baumgardt (2021). Hunt & Reffert (2023) used HDBSCAN algorithm on Gaia DR3 data down to magnitude G 20 mags to recover 7167 clusters, including 134 globular clusters. For GCs, we find that the crossmatch between our catalog and Hunt & Reffert (2023) varies from 15% (for NGC 6656; very close to the Galactic plane) to 95% based on the distance of the clusters from the Galactic plane, with a median crossmatch accuracy of 84%. NGC 104 is not studied by Hunt & Reffert (2023), hence we are unable to make any crossmatch with this cluster. The membership method of Hunt & Reffert (2023) is found to have a bias on whether the cluster is located in a crowded or sparse field, which is not found in the methods of Vasiliev & Baumgardt (2021). In Figure 12, we show CMDs of two GCs (NGC 3201 and NGC 288) to demonstrate the comparison of the cluster membership obtained from our method (blue), Vasiliev & Baumgardt (2021) in orange and Hunt & Reffert (2023) in green. The figure shows a good agreement between all three methods.
@@
We find a better crossmatch with Hunt & Reffert (2023) for OCs. We find a crossmatch of more than 50% for all clusters (except UBC252 with 19%) and with median crossmatch value of 94%. The unmatched sources are mainly fainter than 18 mags in G band and more importantly, all BSSs in our sample are also classified as cluster members in Hunt & Reffert (2023). Dias et al. (2021) used the classic maximum likelihood approach (Dias et al. 2014; Monteiro et al. 2020) on Gaia-DR2 data to study 2000 open clusters for stars brighter than 18 mags in the G band. The comparison shows that the crossmatch fraction varies from 60% to 96% with a median value of 88%. The difference in the crossmatch is mainly found towards the fainter G magnitudes. When cross-matched only with our BSS sample, we find more than 97% are also classified as members in Dias et al. (2021). Cantat-Gaudin et al. (2020a) applied an artificial neural network in Gaia DR2 photometry down to G = 18 mags to estimate the cluster membership of 2000 OCs. Comparing our cluster members with stars brighter than 18 mags in G band, we find that crossmatch fraction varies from cluster to cluster with a range from 40% to 95%. There are mainly two reasons behind the differences in crossmatch fraction: Cantat-Gaudin et al. (2020a) used Gaia DR2 for the membership instead of DR3 with better astrometry and was unable to recover the cluster member properly towards the fainter magnitude limit. For example, there are no cluster member present with G17 mags for the cluster UBC 252, which leads to only 41% of our cluster member. While we consider the match for BSSs, we notice that out of 434 BSSs candidate in our catalog, 409 (94%) are also part of the cluster member in Cantat-Gaudin et al. (2020a). In Figure 13, we present CMDs of two OCs (NGC 188 and berkeley 39) as examples to display the comparison between our cluster membership (open blue box) to those obtained by Hunt & Reffert (2023) marked in orange points, Cantat-Gaudin et al. (2020a) marked in green cross, and Dias et al. (2021) in red points. The cluster memberships are found to be well-matched among the different methods. Jadhav & Subramaniam (2021) also provides an extensive catalog of BSSs candidates in OCs. They used the primary catalog of OCs by Cantat-Gaudin et al. (2020a), however. We therefore did not perform a detailed cross-match with Jadhav & Subramaniam (2021). Therefore, we find that our cluster membership matches well with the literature and almost all BSSs are also classified as members in the literature, which also validates the membership determination algorithm used in our study.